OpenStudy (anonymous):

In a circle, a 90(degree) sector has an area of 36pi in^2. What is the radius of this circle?

OpenStudy (anonymous):

since it is 90 degrees, it is a quarter of the area of the whole scircle area = pi(r^2)/4 36 pi = pi (r^2)/4 pi cancels 144 = r^2 r = 12

OpenStudy (anonymous):

Thank you @jkasper !

OpenStudy (anonymous):

no problem